# An Introduction to Zero-Order Optimization Techniques for Robotics

### v1

@@ IV-C Connection to the Reinforce algorithm
$\frac{\partial}{\partial\theta}\log(p_{\theta}(a|s))=\frac{\partial\pi_{\theta}(s)}{\partial\theta}^{\top}\Sigma^{-1}(a-\pi_{\theta}(s))$ (37)
@@
$\left.\frac{\partial\pi_{\theta}(S_{k})}{\partial\theta}^{\top}\Sigma^{-1}(A_{k}-\pi_{\theta}(S_{k}))\right]$ (38)
The baseline $V(S_{k})$ can be subtracted because the random variable $A_{k}-\pi_{\theta}(S_{k})$ has zero-mean conditioned on the state. Similarly to the randomized smoothing case this induces an invariance to translation by a constant of the reward function. Interestingly, we recover a formula very similar to Equation (34). The only difference is that the rollouts are now stochastic. This changes two things. The first one is that the $Q$-function and the partial derivatives are now evaluated at $S_{k}$ (instead of $s_{k}$ in Equation (34)). The second is that the definition of the $Q$-function has changed. In Equation (34), the $Q$-function denoted deterministic rollouts. In contrast, it is now an expectation of the outcome of stochastic rollouts.
The additional stochasticity in Equation (IV-C) (as compared to Equation (34)) can be understood as exploration noise. One formal justification could be to understand the stochastic policy in Equation (35) as a smoothing operating in action space. Similarly to Gaussian smoothing, the smoothing allows to derive the gradient of the cost with only $Q$-function evaluation. Consequently, the policy gradient theorem is analogous to the log-likelihood trick of Gaussian smoothing.
In the end, Reinforce can be seen as a way to solve the deterministic formulation introduced in Equation (29) without the simulator’s gradients by sampling in the action space. Interestingly, the smoothing of the Reinforce formulation introduces additional stochasticity as compared to Equation (34). It would be interesting to understand whether this extra stochasticity that can be interpreted as exploration noise is useful in practice. Intuitively, smoothing helps to simplify the problem but adding too much noise might render the surrogate too different from the initial problem. Future work could explore connections to other popular stochastic policy gradient algorithms such as TRPO or PPO.
@@ V Population based algorithms
Theoretically, sampling-based algorithms following Langevin dynamics can sample from any probability distribution [23]. However, when applied to challenging multimodal probability distributions, the samples may fail to mix well and can exhibit strong autocorrelation. A natural way to address this issue is to explore algorithms that consider multiple starting points.
So far, all the algorithms that we have seen, either in TO or RL, have been local search techniques based on the simple greedy local search from Algorithm 2. It would be worth deriving algorithms combining local search concepts with global ones.
@@ 6             $x\leftarrow\tilde{x}$
A more sophisticated approach is to perform several local search in parallel. One way to do this is to maintain a population of $N$ samples, generate a new population by applying local search and then select the $N$ best samples. This procedure, called $(N+\lambda)$-ES, is one of the most popular ES algorithms. Algorithm 9 summarizes this procedure. Interestingly, there are two special cases: (1 + 1)-ES is equivalent to Greedy Local Search (Algorithm 2) and (1 + $\lambda$)-ES is equivalent to Predictive sampling [9] (Algorithm 4).
Input: Initial population of $N$ values: $x_{1},\dots,x_{N}$
@@ 6             Add $x+d$ to $D$
8      $\{x_{1},\dots,x_{N}\}\leftarrow$ the $N$ elements in $D$ with smallest $f(x)$
@@
A potential limitation of this type of approach is that many greedy local searches may converge to the same local minima. Ideally, samples should coordinate with each other. An intuitive idea would be to avoid having particles too close to each other. One way to do so is to perform Stein Variational Gradient Descent (SVGD) [41]. This formulation naturally encompasses a repulsion term to maintain a certain distance between particles. Although SVGD requires gradients, it is natural to use the concepts covered in the previous Sections to derive a gradient free variation. For instance, [42] derives a gradient free SVGD based on CMA-ES.
VI Conclusion
@@ Appendix A Connection between the greedy local search and the Metropolis-Hastings algorithm
[26] shows how the Metropolis adjusted Langevin algorithm can be interpreted as a line search with the Armijo condition. Similarly, in the trivial case where the step size $\alpha_{k}=0$, in Equation (14), we can recover a connection to Algorithm 2. The idea of the Metropolis-Hastings algorithm [23] is to accept a step from $x$ to $x^{\prime}$ with probability:
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right)$ (39)
@@
$q(\cdot|x)\propto\mathcal{N}(x,\gamma^{2}I)$ (40)
@@
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right)=\min\left(1,e^{-\left(f(x^{\prime})-f(x)\right)}\right)$
Interestingly, $e^{-\left(f(x^{\prime})-f(x)\right)}>1$ implies $f(x^{\prime})<f(x)$. Hence, one could easily transform the Greedy Local Search (Algorithm 2) in an MCMC method with the Metropolis-Hastings algorithm[23]. Indeed, in Algorithm 2, we could accept the step $x^{\prime}=x+d$ with probability one in case of decrease and allow for an increase with probability $e^{-\left(f(x^{\prime})-f(x)\right)}$.
@@ Appendix B About the natural gradient
$\mathcal{L}(h,\lambda)=f(x)+\nabla f(x)^{T}h-\lambda(\frac{1}{2}h^{T}Ah-\epsilon^{2})$ (43)
@@
$\iff h=\frac{1}{\lambda}A^{-1}\nabla f(x)$ (44)
The sign of $\lambda$ can be recovered with the second order optimality condition [43].
$\nabla^{2}_{hh}\mathcal{L}(h,\lambda)\succeq 0\implies-\lambda A\succeq 0\implies\lambda\leq 0$ (45)
@@
$\operatorname{KL}(p_{\theta+h}||p_{\theta})=\int_{x}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$ (47)
Locally, we can show that this metric is quadratic. Let’s look at the partial derivatives of Equation (47) with respect to $h$.
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}=\int_{x}\partial_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$
[…]
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}(h=0)=0$ (48)
Then, we can compute the Hessian with respect to $h$.
$\frac{\partial^{2}\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial^{2}h}=\int_{x}\partial^{2}_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$
$\;\;\;\;\;\;\;\;+\int_{x}\partial_{h}p_{\theta+h}(x)^{T}\partial_{h}p_{\theta+h}(x)\frac{1}{p_{\theta+h}(x)}dx$ (49)
And, we find that:
@@
$=\mathbb{E}_{X\sim p_{\theta}}\left[\partial_{\theta}\log(p_{\theta}(X))^{T}\partial_{\theta}\log(p_{\theta}(X))\right]=F(\theta)$
@@
$\operatorname{KL}(p_{\theta+h}||p_{\theta})=\frac{1}{2}h^{T}F(\theta)h+o(||h||^{2})$ (51)
@@
$p_{\theta}(x)=\frac{1}{\kappa}\exp\left(-\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)\right)$
@@
$\partial_{\theta}\log(p_{\theta}(x))=(x-\theta)^{T}\Sigma^{-1}$ (52)
@@ Appendix C About the connection between the log-sum-exp and risk-sensitive control
Interestingly, the surrogate defined in Equation (18) reminds of the risk-seeking control formulation [29]. At first, this can seem counter-intuitive as in practice, we prefer designing risk-averse controllers rather than risk-seeking ones. However, from an optimization perspective, the risk-seeking formulation has the advantage of giving more importance to lower values. Figure 4 illustrates the landscape for each formulation. The risk-neutral formulation corresponds to the default randomized smoothing. One way to interpret these landscapes is to see the smoothing noise as a disturbance from the environment. On one hand, the risk-seeking formulation expects the noise to act in its favor by shifting the state towards the global solution. Therefore, it will lower the function value in the surroundings of the narrow global minima. On the other hand, the risk-averse formulation expects the noise to be adversarial. Therefore, it will increase the function value (as compared to the neutral formulation) around the narrow global minima.
@@ Appendix D Connection to the Cross Entropy Method (CEM)
The Cross Entropy Method (CEM) is a popular algorithm that can be used for optimization. Algorithm 10 summarizes the main step of CEM. Interestingly, this algorithm connects to the natural gradient [30]. Indeed, in Equation (III-C), one can assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 00 to the others (by best, we mean samples with the lowest value $f(x_{k})$). Choosing a step-size $\alpha=1$ yields an algorithm close to the well-known Cross-Entropy Method (CEM). The only difference with the original CEM is that natural gradient update computes the covariance update with the old $x$ (i.e. the one used to sample) instead of the new one [30].
@@ Appendix E Additional TO results
Figure 2 provides additional results for the TO benchmark. Specifically, we show how block-diagonal CMA compares to the default CMA (relying on the full matrix). In addition, we illustrate the advantage of having different step-sizes for the mean and the covariance.

### v2

@@ IV-D Connection to the Reinforce algorithm
$\frac{\partial}{\partial\theta}\log(p_{\theta}(a|s))=\frac{\partial\pi_{\theta}(s)}{\partial\theta}^{\top}\Sigma^{-1}(a-\pi_{\theta}(s)).$ (39)
@@
$\left.\frac{\partial\pi_{\theta}(S_{k})}{\partial\theta}^{\top}\Sigma^{-1}(A_{k}-\pi_{\theta}(S_{k}))\right].$ (40)
The baseline $V(S_{k})$ can be subtracted because the random variable $A_{k}-\pi_{\theta}(S_{k})$ has a zero-mean conditioned on the state. Similarly to the randomized smoothing case, this induces invariance to translation by a constant of the reward function. Interestingly, we recover a formula very similar to Equation (36). The only difference is that the rollouts are now stochastic. This changes two things. The first one is that the $Q$-function and the partial derivatives are now evaluated at $S_{k}$ (instead of $s_{k}$ in Equation (36)). The second is that the definition of the $Q$-function has changed. In Equation (36), the $Q$-function denotes deterministic rollouts. In contrast, it is now an expectation of the outcome of stochastic rollouts.
The additional stochasticity in Equation (IV-D) (as compared to Equation (36)) can be understood as exploration noise. One formal justification could be to understand the stochastic policy in Equation (37) as a smoothing operating in action space. Similarly to Gaussian smoothing, the smoothing allows one to derive the gradient of the cost with only $Q$-function evaluation. Consequently, the policy gradient theorem is analogous to the log-likelihood trick of Gaussian smoothing.
In the end, Reinforce can be seen as a way to solve the deterministic formulation introduced in Equation (31) without the simulator’s gradients by sampling in the action space. Interestingly, the smoothing of the Reinforce formulation introduces additional stochasticity as compared to Equation (36). It would be interesting to understand whether this extra stochasticity, which can be interpreted as exploration noise, is useful in practice. Intuitively, smoothing helps to simplify the problem, but adding too much noise might render the surrogate too different from the initial problem. Future work could explore connections to other popular stochastic policy gradient algorithms, such as TRPO or PPO.
@@ V Population based algorithms
Theoretically, sampling-based algorithms based on Langevin dynamics can generate samples from any target probability distribution [27]. However, when applied to challenging multimodal distributions, the samples may fail to mix well and can exhibit strong autocorrelation. A natural approach to mitigate this issue is to explore algorithms that consider multiple starting points.
Thus far, all the algorithms that we have considered, either in TO or RL, have been local search techniques based on the simple greedy local search from Algorithm 2. It would be worth deriving algorithms that combine local search concepts with global ones, in the spirit of Algorithm 1.
@@ 6    $x\leftarrow\tilde{x}$
A more sophisticated approach is to perform several local searches in parallel. One way to do this is to maintain a population of $N$ samples, generate a new population by applying local search, and then select the $N$ best samples. This procedure, called $(N+\lambda)$-ES, is one of the most popular Evolutionary Strategy (ES) algorithms. Algorithm 9 summarizes this procedure. Interestingly, there are two special cases: (1 + 1)-ES is equivalent to Greedy Local Search (Algorithm 2) and (1 + $\lambda$)-ES is equivalent to Predictive sampling [9] (Algorithm 4).
Input: Initial population of NN values: $x_{1},\dots,x_{N}$
@@ 6      Add $x+d$ to $D$
8 $\{x_{1},\dots,x_{N}\}\leftarrow$ the NN elements in $D$ with smallest $f(x)$
@@
A potential limitation of this type of approach is that many greedy local searches may converge to the same local minimum. Ideally, samples should coordinate with one another. An intuitive idea is to avoid having particles that are too close to one another. One way to do so is to perform Stein Variational Gradient Descent (SVGD) [44]. This formulation naturally encompasses a repulsion term to maintain a certain distance between particles. Although SVGD requires gradients, it is natural to use the concepts covered in the previous Sections to derive a gradient-free variation. For instance, [45] derive a gradient-free SVGD based on CMA-ES.
VI Parallel computing
A common feature of zero-order TO and RL is that both require numerous function evaluations. To improve computational efficiency, a typical strategy is to parallelize these evaluations. In the machine learning community, parallel computing is well established with libraries such as JAX [46] and PyTorch [47]. However, in robotics, these libraries cannot be used directly, as each function evaluation relies on a simulator. Indeed, the function evaluation in Equation (17) or Equation (31) involves propagating the system dynamics forward in time. For this reason, parallel simulators such as MuJoCo XLA (MJX) [48] and Isaac Sim [49] have been developed to support massively parallel rollouts on GPUs, which makes them especially suitable for sampling-based TO and RL.
On the algorithmic side, Evosax [50] provides implementations of random search and popular ES algorithms based on JAX. These can then be used for zero-order MPC with Hydrax [37]. Finally, CleanRL [43] offers simple and efficient implementations of widely used RL algorithms.
VII Conclusion
@@
Acknowledgments
We sincerely thank Frederike Dümbgen and Olivier Sigaud for their constructive feedback, which greatly improved this paper.
@@ Appendix A Connection between the greedy local search and the Metropolis-Hastings algorithm
[30] shows how the Metropolis adjusted Langevin algorithm can be interpreted as a line search with the Armijo condition. Similarly, in the trivial case where the step size $\alpha_{k}=0$ in Equation (16), we can recover a connection to Algorithm 2. The idea of the Metropolis-Hastings algorithm [27] is to accept a step from $x$ to $x^{\prime}$ with probability:
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right).$ (41)
@@
$q(\cdot|x)\propto\mathcal{N}(x,\gamma^{2}I).$ (42)
@@
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right)=\min\left(1,e^{-\left(f(x^{\prime})-f(x)\right)}\right).$
Interestingly, $e^{-\left(f(x^{\prime})-f(x)\right)}>1$ implies $f(x^{\prime})<f(x)$. Hence, one could easily transform the Greedy Local Search (Algorithm 2) into an MCMC method with the Metropolis-Hastings algorithm [27]. Indeed, in Algorithm 2, we could accept the step $x^{\prime}=x+d$ with a probability of one in the case of a decrease and allow for an increase with a probability of $e^{-\left(f(x^{\prime})-f(x)\right)}$.
@@ Appendix B About the natural gradient
$\mathcal{L}(h,\lambda)=f(x)+\nabla f(x)^{T}h-\lambda(\frac{1}{2}h^{T}Ah-\epsilon^{2}).$ (45)
@@
$\iff h=\frac{1}{\lambda}A^{-1}\nabla f(x).$ (46)
The sign of $\lambda$ can be recovered with the second order optimality condition [51]:
$\nabla^{2}_{hh}\mathcal{L}(h,\lambda)\succeq 0\implies-\lambda A\succeq 0\implies\lambda\leq 0.$ (47)
@@
$\hskip-5.69046pt\operatorname{KL}(p_{\theta+h}||p_{\theta})=\int_{x}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx.$ (49)
Locally, we can show that this metric is quadratic. Let’s look at the partial derivatives of Equation (49) with respect to $h$:
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}=\int_{x}\partial_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx.$
As $\int_{x}p_{\theta+h}(x)dx=1$ implies $\partial_{h}\int_{x}p_{\theta+h}(x)dx=0$ for all $h$. Hence,
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}(h=0)=0.$ (50)
Then, we can compute the Hessian with respect to $h$:
$\hskip-8.5359pt\frac{\partial^{2}\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial^{2}h}=\int_{x}\partial^{2}_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$
$\;\;\;\;\;\;\;\;+\int_{x}\partial_{h}p_{\theta+h}(x)^{T}\partial_{h}p_{\theta+h}(x)\frac{1}{p_{\theta+h}(x)}dx.$ (51)
And we find that:
@@
$=\mathbb{E}_{X\sim p_{\theta}}\left[\partial_{\theta}\log(p_{\theta}(X))^{T}\partial_{\theta}\log(p_{\theta}(X))\right]=F(\theta),$
@@
$\operatorname{KL}(p_{\theta+h}||p_{\theta})=\frac{1}{2}h^{T}F(\theta)h+o(||h||^{2}).$ (53)
@@
$p_{\theta}(x)=\frac{1}{\kappa}\exp\left(-\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)\right).$
@@
$\partial_{\theta}\log(p_{\theta}(x))=(x-\theta)^{T}\Sigma^{-1}.$ (54)
@@ Appendix C About the connection between the log-sum-exp and risk-sensitive control
Interestingly, the surrogate defined in Equation (20) resembles the risk-seeking control formulation [33]. At first, this may seem counter-intuitive, as in practice, we prefer designing risk-averse controllers rather than risk-seeking ones. However, from an optimization perspective, the risk-seeking formulation has the advantage of giving more importance to lower values. Figure 4 illustrates the landscape for each formulation. The risk-neutral formulation corresponds to the default randomized smoothing. One way to interpret these landscapes is to see the smoothing noise as a disturbance from the environment. On one hand, the risk-seeking formulation expects the noise to act in its favor by shifting the state towards the global solution. Therefore, it lowers the function value in the vicinity of the narrow global minima. On the other hand, the risk-averse formulation expects the noise to be adversarial. Therefore, it increases the function value (as compared to the neutral formulation) around the narrow global minima.
@@ Appendix D Connection to the Cross Entropy Method (CEM)
The Cross Entropy Method (CEM) is a popular algorithm that can be used for optimization. Algorithm 10 summarizes the main steps of CEM. Interestingly, this algorithm connects to the natural gradient [34]. Indeed, in Equation (III-C), one can assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 0 to the others (by best, we mean samples with the lowest value $f(x_{k})$). Choosing a step-size $\alpha=1$ yields an algorithm close to the well-known Cross-Entropy Method (CEM). The only difference with the original CEM is that the natural gradient update computes the covariance update using the old $x$ (i.e. the one used to sample) instead of the new one [34].
@@ Appendix E Additional TO results
Figure 2 provides additional results for the TO benchmark. Specifically, we show how block-diagonal CMA compares to the default CMA (which relies on the full matrix). In addition, we illustrate the advantage of having different step-sizes for the mean and the covariance.